OpenStudy (anonymous):

Find the length of the missing leg of a right triangle given a leg of length 8 and a hypotenuse of length 10. Leave your answer in simplest radical form. (1 point)

OpenStudy (raden):

let the length of the missing side is x, then use the pythagorean theorem : x^2 = 10^2 - 8^2 solve for x

OpenStudy (anonymous):

would it be d?

OpenStudy (raden):

just guess ?

OpenStudy (anonymous):

no. i am failing geometry!

OpenStudy (raden):

well, to get x we have to simply the equation above : x^2 = 10^2 - 8^2 x^2 = 100 - 64 x^2 = 36 x = sqrt (36) = ............... ?

OpenStudy (anonymous):

so 6.

OpenStudy (raden):

yep

OpenStudy (anonymous):

k help me with one more?

OpenStudy (raden):

ok

OpenStudy (anonymous):

Find the length of the missing side. Leave your answer in simplest radical form. The triangle is not to scale. 25 144 5

OpenStudy (raden):

diagram ?

OpenStudy (anonymous):

d is square root of 5.

OpenStudy (anonymous):

yes. and d is square root of 5.

OpenStudy (raden):

let the length of hypotenusa is x, and use the pythagorean theorema again : x^2 = 3^2 + 4^2 x^2 = 9 + 16 x^2 = 25 x = sqrt(25) = .......................................................... ?

OpenStudy (anonymous):

so 5 just 5 no square root of 5? just simply 5?

OpenStudy (raden):

5 is the answer

OpenStudy (raden):

sqrt(25) = 5